# On the Existence of Cyclic Lattice Codes

### v1

@@ II-C Rectangular Encoding
$Q_{\Lambda_{s}}(\mathbf{y})=\min_{\mathbf{x}\in\Lambda_{s}}\left\|\mathbf{y}-\mathbf{x}\right\|^{2}.$ (6)
@@ II-D Cyclic Groups, Cyclic Lattice Codes
A nested lattice code $\mathcal{C}$ is called a cyclic lattice code [^1] In this paper, the term ‘cyclic lattice code’ refers to a cyclic group and in general they are not cyclic codes which is a block code where a cyclic shift of each codeword also gives a codeword. in coordinate $k$ if it is a cyclic group of order $M$ and can be generated by $\mathbf{g}_{k}$, which is the column $k$ of $\mathbf{G}_{c}$. In other words, the whole codebook $\mathcal{C}$ can be generated from information vector $\mathbf{b}=\begin{bmatrix}0&\dots&0&b_{k}&0&\dots&0\end{bmatrix}^{T}$ where $b_{k}\in\left\{0,1,...,M-1\right\}$ with operation
@@
where $\mathbf{y}\in\mathcal{C}$ and $\mathbf{g}_{k}$ is the column $k$ of generator matrix $\mathbf{G}_{c}$. In this case, the encoding matrix $\mathbf{M}$ satisfies $M_{k}=M$ and $M_{i}=1$ for any $i\neq k$ with $i,k\in\{1,\dots,n\}$ [2].
@@ III Existence of Cyclic Lattices
Lemma 1 is a preliminary step needed for proving the central result Lemma 2.
Lemma 1: Consider an $n$-dimension lattice $\Lambda_{s}$ with generator matrix $\mathbf{G}_{s}=\begin{bmatrix}\mathbf{g}_{1}&\mathbf{g}_{2}&\dots&\mathbf{g}_{n}\end{bmatrix}$. The line segment with endpoint $\mathbf{0}$ and $\mathbf{y}=\mathbf{G}_{s}\cdot\mathbf{b}$ with $\mathbf{b}=\begin{bmatrix}b_{1}&b_{2}&\dots&b_{n}\end{bmatrix}^{T}$ does not intersect any other point of $\Lambda_{s}$ if and only if $\gcd(b_{1},b_{2},...,b_{n})=1$.
@@
Consider a cyclic lattice code $\mathcal{C}$ with sublattice condition $\Lambda_{s}\subseteq\Lambda_{c}$. Without loss of generality, we choose $\mathbf{g}_{n}$, column $n$ of $\mathbf{G}_{c}$, as the generator of the cyclic group. Because the group generator is $\mathbf{g}_{n}$, $M\cdot\mathbf{g}_{n}\bmod\Lambda_{s}=\mathbf{0}$ holds, and so it must be that
@@
$\mathcal{P}(\mathbf{G}_{c}\mathbf{M})$ is fundamental region of $\Lambda_{s}$ if and only if $k\cdot\mathbf{g}_{n}\bmod\Lambda_{s}\neq\mathbf{0}$ for $k=1,\dots,M-1$. Following Lemma 1, we could therefore see that $\mathcal{C}$ is cyclic if and only if $\gcd(s_{1},\ldots,s_{n})=1$. Since $\mathbf{H}_{s}\mathbf{G}_{c}=(\mathbf{H}_{c}\mathbf{G}_{s})^{-1}$, (10) could be written as
@@
where $\mathbf{q}_{i}$ is the column $i$ of $\text{det}(\mathbf{H}_{c}\mathbf{G}_{s})(\mathbf{H}_{c}\mathbf{G}_{s})^{-1}$. Since $\text{det}(\mathbf{H}_{c}\mathbf{G}_{s})=M$, we have
@@
Lemma 2: An $n$ dimensional nested lattice code $\mathcal{C}$ with $\Lambda_{s}\subseteq\Lambda_{c}$ is a cyclic lattice code in coordinate $i\in\left\{1,2,...,n\right\}$ if and only if $\gcd(\mathbf{q}_{i})=1$, where $\mathbf{q}_{i}$ is the column $i$ of $\text{det}(\mathbf{H}_{c}\mathbf{G}_{s})(\mathbf{H}_{c}\mathbf{G}_{s})^{-1}$.
@@ III-A $n=2$ Nested Lattice
First, we would like to design an $n=2$ cyclic lattice code using Lemma 2. Suppose we have
@@
and $\mathbf{g}_{2}$ of $\mathbf{G}_{c}$ is chosen as generator, then (12) becomes
$\begin{bmatrix}-b&a\end{bmatrix}=\begin{bmatrix}s_{1}&s_{2}\end{bmatrix}$ (14)
@@
Since pair $(4,9)$ and $(3,8)$ are both relatively prime, either coordinate $2$ or $1$ can be chosen respectively, i.e., $\mathbf{g}_{2}=\begin{bmatrix}\frac{2}{9}&\frac{8}{9}\end{bmatrix}^{T}$ or $\mathbf{g}_{1}=\begin{bmatrix}\frac{4}{3}&\frac{4}{3}\end{bmatrix}^{T}$ can be generators. In this case, size of $\mathcal{C}$ is $M=5$ and therefore we could have either $\mathbf{M}=\mathrm{diag}(1,5)$ for coordinate $2$ or $\mathbf{M}=\mathrm{diag}(5,1)$ for coordinate $1$. Applying (5), we can obain the codebook $\mathcal{C}$, which is shown in Fig. 1.
@@ III-B Design for General $n$
For general $n\geq 3$, there are more degrees of freedom to design $\mathbf{H}_{c}\mathbf{G}_{s}$ using Lemma 2. We would like to restrict $\mathbf{H}_{c}\mathbf{G}_{s}$ to make the design easier. Denote $\mathbf{W}=\mathbf{H}_{c}\mathbf{G}_{s}$ and
@@
A sufficient condition to have $\gcd(s_{1},s_{2},...,s_{n})=1$, is that a subset $\mathcal{S}^{\prime}\subseteq\left\{s_{1},s_{2},...,s_{n}\right\}$ has $\gcd(\mathcal{S}^{\prime})=1$, where $2\leq|\mathcal{S}^{\prime}|\leq n$. We give simple designs of $\mathbf{W}$ matrix for $|\mathcal{S}^{\prime}|=2$ and $|\mathcal{S}^{\prime}|=3$.
@@
where $\gcd(a,b)=1$. The first row except $a,b$ are all zeros, the submatrix start from row $2$ to row $n-1$, column $1$ to column $n-2$ is a unimodular matrix which we simply choose as the identity matrix, the remaining values are free to choose as long as $\mathbf{W}$ remains to be full rank. With this special structure, it is trivial to compute $C^{(n,n-1)}=-\text{det}(\mathbf{W}^{(n,n-1)})=-b$ and $C^{(n,n)}=\text{det}(\mathbf{W}^{(n,n)})=a$. Because $a,b$ are relatively prime, $\gcd(C^{(n,n-1)},C^{(n,n)})=1$, satisfying the conditions of Lemma 2.
@@
Thus, if $\gcd(-b+c,a)=1$ then Lemma 2 holds.
@@ IV Conditions for Isomorphism Existence
Nazar and Gastpar [3] proposed the compute-and-forward techniques for physical-layer network coding, where the group properties of lattice codes are essential. Linear labeling of $\Lambda_{c}$ with group linearity properties was proposed by Feng et al. [7]. They also gave the conditions on the existence of an isomorphism. In addition, there are conditions for group isomorphism based on the structure of the matrix $\mathbf{H}_{c}\mathbf{G}_{s}$ [2], which is given in Lemma 3 below.
Define group *$\mathbb{Z}_{*}^{n}$ as:
*$\mathbb{Z}_{*}^{n}=\mathbb{Z}_{M_{1}}\times\mathbb{Z}_{M_{2}}\times\dots\times\mathbb{Z}_{M_{n}},$ (22)
where $\mathbb{Z}_{M_{i}}=\mathbb{Z}/M_{i}\mathbb{Z}=\left\{0,1,...,M_{i}-1\right\}$. Consider two information vectors $\mathbf{b}_{1},\mathbf{b}_{2}\in\mathbb{Z}_{*}^{n}$, a group isomorphism is satisfied if
@@
Lemma 3 [2]: For arbitrary nested lattice $\Lambda_{s}\subseteq\Lambda_{c}$, if all elements from row $i$ of $\mathbf{H}_{c}\mathbf{G}_{s}$ could be divided by $M_{i}$ for all $i=1,2,...,n$, then an isomorphism exists between group *$\mathbb{Z}_{*}^{n}$ and $\mathcal{C}$ under group operation $\oplus$.
Clearly, $\mathbf{H}_{c}\mathbf{G}_{s}$ in (16) does not satisfy the condition in Lemma 3 under cyclic encodings since neither row is divisible by $M=5$. On the other hand, a self-similar lattice $\Lambda/K\Lambda$ which could be used as shaping lattice for compute-and-forward relaying [8] satisfies the isomorphism since its rectangular coding matrix $\mathbf{M}=K\mathbf{I}_{n}$ but it is not cyclic.
@@
Eqn. (25) is a linear Diophantine equation in the variable $r_{1},r_{2},...,r_{n}$, which might have no solutions or multiple solutions. For $\mathbf{W}$ in (19), setting $r_{n-1},r_{n}\neq 0$ and the other $r_{i}=0$ gives rise to the linear Diophantine equation
@@
(26) and (27) are naturally solvable as long as the cyclic conditions are satisfied, i.e., $\gcd(a,b)=1$ and $\gcd(-b+c,a)=1$ holds, respectively [9]. Hence, we can come to Lemma 4.
Lemma 4: For a cyclic lattice code of arbitrary size $M$ with structure $\mathbf{W}$ in (19) or (20), it is always possible to keep isomorphism by modifying the last row of $\mathbf{W}$ to the form of (24), which is equivalent to finding a solution to (26) and (27) respectively and setting any other irrelevant $r_{i}$ to 00.
@@ IV-A $n=2$ nested lattice
For the case of $n=2$, $\mathbf{H}_{c}\mathbf{G}_{s}$ only have last two columns of (19). Consider $\mathbf{M}=\mathrm{diag}(1,M)$ and
@@
Since $A_{2}$ lattice has the best shape in 2 dimensions, we also form a nested lattice code of arbitrary size $M$ using shaping lattice $A_{2}$ with generator matrix given by
@@
Using the same $(a,b,r_{1},r_{2})=(4,9,-1,-2)$ as in the last example, and setting $M=11$, we could obtained the nested lattice code as shown in Fig 3.
@@ IV-B $E_{8}$ lattice
Consider an $E_{8}$ lattice as shaping lattice and design $\mathbf{W}$ with structure in (20) where $a=7,b=17,c=19$. One possible solution to (27) is $r_{6}=95,r_{7}=65,r_{8}=92$ and $r_{i}=0,i\leq 5$. Setting $M=64$, we can obtain an isomorphic cyclic lattice code $\mathcal{C}$ with a coding gain of $2.67$ dB, where the coding lattice $\mathbf{G}_{c}=\mathbf{G}_{s}\mathbf{W}^{-1}$. Since we use $E_{8}$ as shaping lattice, the shaping gain is unchanged, i.e., $0.65$ dB. $\mathcal{C}$ is in the subspace $\mathbb{R}^{3}$ which is shown in Fig. 4.
@@ V Conclusion
In this paper, we derived conditions for the existence of cyclic lattice codes and proposed two simple designs for general $n$ dimensional cyclic lattice codes. By solving linear Diophantine equations, these designs can be easily adapted to possess group isomorphism, which makes cyclic lattice code a potential candidate for physical layer network coding. By fixing either the shaping lattice or coding lattice, we can design an isomorphic cyclic lattice code of arbitrary codebook size $M$. Cyclic lattice codes have shaping gain when a suitable shaping lattice is used. In this case, however, the characteristics of the coding gain are still unclear and remains as future work. One might also consider applying this work to physical layer network relaying techniques such as compute and forward [3].

### v2

@@ II-C Rectangular Encoding
$Q_{\Lambda_{s}}(\mathbf{y})=\operatorname*{argmin}_{\mathbf{x}\in\Lambda_{s}}\left\|\mathbf{y}-\mathbf{x}\right\|^{2}.$ (6)
@@ II-D Cyclic Groups, Cyclic Lattice Codes
A nested lattice code $\mathcal{C}$ is called a cyclic lattice code [^1] In this paper, the term “cyclic lattice code” refers to a cyclic group and in general they are not cyclic codes which is a block code where a cyclic shift of each codeword also gives a codeword. Micciancio studied cyclic lattices which are defined analogously to cyclic codes [7]. in coordinate $k$ if it is a cyclic group of order $M$ and can be generated by $\mathbf{g}_{k}$, which is the column $k$ of $\mathbf{G}_{c}$. In other words, the whole codebook $\mathcal{C}$ can be generated from information vector $\mathbf{b}=\begin{bmatrix}0&\dots&0&b_{k}&0&\dots&0\end{bmatrix}^{T}$ where $b_{k}\in\left\{0,1,...,M-1\right\}$ with operation
@@
where $\mathbf{y}\in\mathcal{C}$. In this case, the encoding matrix $\mathbf{M}$ satisfies $M_{k}=M$ and $M_{i}=1$ for any $i\neq k$ with $i,k\in\{1,\dots,n\}$ [2].
@@ III Existence of Cyclic Lattices
Lemma 1 is a preliminary step needed for proving the central result Proposition 2.
Lemma 1: Consider an $n$-dimensional lattice $\Lambda_{s}$ with generator matrix $\mathbf{G}_{s}=\begin{bmatrix}\mathbf{g}_{1}&\mathbf{g}_{2}&\dots&\mathbf{g}_{n}\end{bmatrix}$. The line segment with endpoint $\mathbf{0}$ and $\mathbf{y}=\mathbf{G}_{s}\cdot\mathbf{b}$ with $\mathbf{b}=\begin{bmatrix}b_{1}&b_{2}&\dots&b_{n}\end{bmatrix}^{T}$ does not intersect any other point of $\Lambda_{s}$ if and only if $\gcd(b_{1},b_{2},...,b_{n})=1$.
@@
Consider a cyclic lattice code $\mathcal{C}$ with $\Lambda_{s}\subseteq\Lambda_{c}$. Without loss of generality, we choose $\mathbf{g}_{n}$, column $n$ of $\mathbf{G}_{c}$, as the generator of the cyclic group. Because the group generator is $\mathbf{g}_{n}$, $M\cdot\mathbf{g}_{n}\bmod\Lambda_{s}=\mathbf{0}$ holds, and so it must be that
@@
$\mathcal{P}(\mathbf{G}_{c}\mathbf{M})$ is a fundamental region of $\Lambda_{s}$ if and only if $k\cdot\mathbf{g}_{n}\bmod\Lambda_{s}\neq\mathbf{0}$ for $k=1,\dots,M-1$. Following Lemma 1, we could therefore see that $\mathcal{C}$ is cyclic if and only if $\gcd(s_{1},\ldots,s_{n})=1$. Since $\mathbf{H}_{s}\mathbf{G}_{c}=(\mathbf{H}_{c}\mathbf{G}_{s})^{-1}$, (10) could be written as
@@
where $\mathbf{q}_{i}$ is column $i$ of $\text{det}(\mathbf{H}_{c}\mathbf{G}_{s})(\mathbf{H}_{c}\mathbf{G}_{s})^{-1}$. Since $\text{det}(\mathbf{H}_{c}\mathbf{G}_{s})=M$, we have
@@
Proposition 2: An $n$ dimensional nested lattice code $\mathcal{C}$ with $\Lambda_{s}\subseteq\Lambda_{c}$ is a cyclic lattice code in coordinate $i\in\left\{1,2,...,n\right\}$ if and only if $\gcd(\mathbf{q}_{i})=1$, where $\mathbf{q}_{i}$ is column $i$ of $\text{det}(\mathbf{H}_{c}\mathbf{G}_{s})(\mathbf{H}_{c}\mathbf{G}_{s})^{-1}$.
@@ III-A $n=2$ Nested Lattice
First, we would like to design an $n=2$ cyclic lattice code using Proposition 2. Suppose we have
@@
and $\mathbf{g}_{2}$ of $\mathbf{G}_{c}$ is chosen as the generator of the cyclic group, then (12) becomes
$\begin{bmatrix}-b&a\end{bmatrix}=\begin{bmatrix}s_{1}&s_{2}\end{bmatrix}.$ (14)
@@
Since both pairs $(4,9)$ and $(3,8)$ are both relatively prime, either coordinate $2$ or $1$ can be chosen respectively, i.e., $\mathbf{g}_{2}=\begin{bmatrix}\frac{2}{9}&\frac{8}{9}\end{bmatrix}^{T}$ or $\mathbf{g}_{1}=\begin{bmatrix}\frac{4}{3}&\frac{4}{3}\end{bmatrix}^{T}$ can be generators. The size of $\mathcal{C}$ is $M=5$ and therefore we could have either $\mathbf{M}=\mathrm{diag}(1,5)$ for coordinate $2$ or $\mathbf{M}=\mathrm{diag}(5,1)$ for coordinate $1$. Applying (5), we can obtain the codebook $\mathcal{C}$, which are shown in Fig. 1.
@@ III-B Design for General $n$
For general $n\geq 3$, there are more degrees of freedom to design $\mathbf{H}_{c}\mathbf{G}_{s}$ using Proposition 2. We would like to restrict $\mathbf{H}_{c}\mathbf{G}_{s}$ to make the design easier. Denote $\mathbf{W}=\mathbf{H}_{c}\mathbf{G}_{s}$ and
@@
A sufficient condition to have $\gcd(s_{1},s_{2},...,s_{n})=1$ is that any subset $\mathcal{S}^{\prime}\subseteq\left\{s_{1},s_{2},...,s_{n}\right\}$ has $\gcd(\mathcal{S}^{\prime})=1$, where $2\leq|\mathcal{S}^{\prime}|\leq n$. We give simple designs of the $\mathbf{W}$ matrix for $|\mathcal{S}^{\prime}|=2$ and $|\mathcal{S}^{\prime}|=3$.
@@
where $\gcd(a,b)=1$. The first row except $a,b$ are all zeros, the submatrix from row $2$ to row $n-1$, column $1$ to column $n-2$ is a unimodular matrix which we simply choose as the identity matrix, the remaining values are free to choose as long as $\mathbf{W}$ remains to be full rank. With this special structure, it is trivial to compute $C^{(n,n-1)}=-\text{det}(\mathbf{W}^{(n,n-1)})=-b$ and $C^{(n,n)}=\text{det}(\mathbf{W}^{(n,n)})=a$. Because $a,b$ are relatively prime, $\gcd(C^{(n,n-1)},C^{(n,n)})=1$, satisfying the conditions of Proposition 2.
@@
Thus, if $\gcd(-b+c,a)=1$ then Proposition 2 holds.
@@ IV Conditions for Isomorphism Existence
Nazar and Gastpar [3] proposed compute-and-forward (CF) techniques for physical-layer network coding. Several algebraic approaches have been proposed for constructing nested lattice codes which the CF framework relies on, e.g., [8, 9, 10, 11, 12]. In general, the group properties of nested lattice codes are essential. Linear labeling of $\Lambda_{c}$ with group linearity properties was proposed by Feng et al. [8]. They also gave conditions on the existence of an isomorphism. In addition, there are conditions for group isomorphism based on the structure of the matrix $\mathbf{H}_{c}\mathbf{G}_{s}$ [2], which is given in Lemma 3 below.
Define group $\mathbb{Z}_{*}^{n}$ as:
$\mathbb{Z}_{*}^{n}=\mathbb{Z}_{M_{1}}\times\mathbb{Z}_{M_{2}}\times\dots\times\mathbb{Z}_{M_{n}},$ (22)
where $\mathbb{Z}_{M_{i}}=\mathbb{Z}/M_{i}\mathbb{Z}=\left\{0,1,...,M_{i}-1\right\}$. Consider two information vectors $\mathbf{b}_{1},\mathbf{b}_{2}\in\mathbb{Z}_{*}^{n}$. A group isomorphism is satisfied if
@@
Lemma 3 [2]: For arbitrary nested lattice $\Lambda_{s}\subseteq\Lambda_{c}$, if all elements from row $i$ of $\mathbf{H}_{c}\mathbf{G}_{s}$ could be divided by $M_{i}$ for all $i=1,2,...,n$, then an isomorphism exists between group $\mathbb{Z}_{*}^{n}$ and $\mathcal{C}$ under group operation $\oplus$.
Clearly, $\mathbf{H}_{c}\mathbf{G}_{s}$ in (16) does not satisfy the condition in Lemma 3 under cyclic encodings since neither row is divisible by $M=5$. On the other hand, a self-similar lattice code $\Lambda/K\Lambda$ which could be used for compute-and-forward relaying [13] satisfies the isomorphism since its rectangular coding matrix $\mathbf{M}=K\mathbf{I}_{n}$ but it is not cyclic.
@@
Eqn. (25) is a linear Diophantine equation in the variables $r_{1},r_{2},...,r_{n}$, which might have no solutions or multiple solutions. For $\mathbf{W}$ in (19), setting $r_{n-1},r_{n}\neq 0$ and the other $r_{i}=0$ gives rise to the linear Diophantine equation
@@
(26) and (27) are naturally solvable as long as the cyclic conditions are satisfied, i.e., $\gcd(a,b)=1$ and $\gcd(-b+c,a)=1$ holds, respectively. Hence, we can come to Lemma 4.
Lemma 4: For a cyclic lattice code of arbitrary size $M$ with structure $\mathbf{W}$ in (19) or (20), there always exists an isomorphsim that can be obtained by modifying the last row of $\mathbf{W}$ to the form of (24), which is equivalent to finding a solution to (26) and (27) respectively and setting any other irrelevant $r_{i}$ to 00.
@@ IV-A $n=2$ nested lattice
For the case of $n=2$, $\mathbf{H}_{c}\mathbf{G}_{s}$ only has the last two columns of (19). Consider $\mathbf{M}=\mathrm{diag}(1,M)$ and
@@
Since the $A_{2}$ lattice has the best shape in 2 dimensions, we also form a nested lattice code of arbitrary size $M$ using shaping lattice $A_{2}$ with generator matrix given by
@@
Using the same $(a,b,r_{1},r_{2})=(4,9,-1,-2)$ as in the last example, and setting $M=11$, we could obtain the isomorphic and cyclic nested lattice code as shown in Fig 3.
@@ IV-B $E_{8}$ lattice
Consider the $E_{8}$ lattice as a shaping lattice and design $\mathbf{W}$ with structure in (20) where $a=7,b=17,c=19$. One possible solution to (27) is $r_{6}=95,r_{7}=65,r_{8}=92$ and $r_{i}=0,i\leq 5$. Setting $M=64$, we can obtain an isomorphic cyclic lattice code $\mathcal{C}$, where the coding lattice $\mathbf{G}_{c}=\mathbf{G}_{s}\mathbf{W}^{-1}$. Since we use $E_{8}$ as shaping lattice, the shaping gain is unchanged, i.e., $0.65$ dB. $\mathcal{C}$ is in the subspace $\mathbb{R}^{3}$ which is shown in Fig. 4.
@@ V Conclusion
In this paper, we derived conditions for the existence of cyclic lattice codes and proposed two simple designs for general $n$ dimensional cyclic lattice codes. By solving linear Diophantine equations, these designs can be easily adapted to possess group isomorphism, which makes cyclic lattice codes a potential candidate for physical layer network coding. By fixing either the shaping lattice or coding lattice, we can design an isomorphic cyclic lattice code of arbitrary codebook size $M$. Cyclic lattice codes have shaping gain when a suitable shaping lattice is used. In this case, however, it might be time consuming to find a good coding lattice. The trade-off between coding gain and shaping gain is still unclear under this design. Nevertheless, it is possible to evaluate the performance with other constructions, e.g., [8, 9, 10, 11, 12] for physical layer network relaying techniques such as compute and forward [3], which remains to be our future work.